Notes of Muon Cooling Experiment Meeting Present: E Black, D. Kaplan, A Moretti, J.Norem, A. Sery, N. Holtkamp, P. Lebrun, D. Carey, D Neuffer. J. Corlett, Steve Kahn (phonelink) Date: Friday, Jan. 15th Notes: Norbert Holtkamp Top 1: Al Moretti gave a status of the work going on in Lab G. The two large magnets which are still in Lab G are going to be removed next week. Beams division so far has not worked on cleaning up the area and technicians are supposed to go out there soon. The Equippment from the PET experiment is still there and will partially be reused. Two Layouts have been worked out so far to accommodate the RF equippment and the magnet+cryogenics. A final decision on the one to be taken has not been made. Schielding blocks will be placed all around the building. For the roof, nothing is foreseen yet. The first experiments will produce only a few MeV beams (max) and radiation should not be an issue with the shielding planned so far. For the open cell rf structure Al presented a 5 cell cavity with the iris matching the beam size. A Shunt Impedance (no transit) Z_o = 25 MOhm/m has been achieved. Ep/E_o is 2.8. A 60 MW power source is required to achieve the appropriate gradients. AL Moretti and Norbert Holtkamp are going to visit a number of companies in the Bay area to discuss the issues of developing high power sources at frequencies between 400 - 800 MHz. Al and John Corlett gave an update on the windows for the beryllium cavity. The windows are supposed to come in friday afternoon (at LBL). John will keep 2, 1 is going to FERMI. They will be used in the low power test of the cavity. The machining for this cavity is not done. At least two more weeks are required to finish. The cavity will not be at LBNL before end of January. Top 2: Ed Black presented a preliminary schedule for the construction and set-up of the high field alternating solenoids and the cooling cavity. The schedule will be further worked out next week with the experts. John Miller from the high field lab will be in Chicago at Feb 22nd (Monday and Tuesday). The developments and schedule for this year will be discussed. Questions which have to be raised are: Do we have enough SSC conductor and can we use it ?? Which are the dimensions for the solenoid?? (That will depend on the cavity to be build (side coupled or not) . Top 3: Steve Kahn gave a quick update on the TPC. He is in the middle of doing more calculations on the resolution and will give more details soon. Top 4: Dan Kaplan presented a number of calculations he has done using dpgeant. He especially optimized the taper of the solenoidal field along the cooling channel to match the smaller emittance and to increase the cooling factor. By almost doubling the field (15 -27..28 T) he increased the cooling factor by 2. The transmission does not significantly change (within the noise). Using the optimizer to further improve the situation turned out to be not useful at this point. Help from Linear optics matching is required and Wishi Wan will be asked to do that. The problems using the optimizer at this point are obvious: Too long computation time or too large errors in the statistics. Dan also looked into what happened in the first two sections where the emittance first seems to grow before it starts shrinking. No Conclusive answer could be given at this point. He looked into the development of the correlation and only could see that a correlation was coming in, which it is supposed to be.